A jet aircraft is traveling at 230 m/s in horizontal flight. The engine takes in air at a rate of 96.7 kg/s and burns fuel at a rate of 3.96 kg/s. The exhaust gases are ejected at 482 m/s relative to the aircraft. Find the thrust of the jet engine. Answer in units of N.


1
Expert's answer
2021-12-12T16:43:46-0500

"F_N=(m_{air}+m_{fuel})V_e-m_{air}V"

where "m_{air}" - the mass rate of air flow through the engine,

"m_{fuel}" - the mass rate of propellant flow entering the engine

"V_e" - the speed of the exhaust gases relative to the aircraft

V - the velocity of the air intake

"F_N=(96.7+3.96)\times482-96.7\times 230=48518.12-22241=26277.12N"